### v1

# FLAME[Uncaptioned image]: Frozen Large Language Models Enable Data-Efficient Language-Image Pre-training

Abstract
Language-image pre-training faces significant challenges due to limited data in specific formats and the constrained capacities of text encoders. While prevailing methods attempt to address these issues through data augmentation and architecture modifications, they continue to struggle with processing long-form text inputs, and the inherent limitations of traditional CLIP text encoders lead to suboptimal downstream generalization. In this paper, we propose FLAME (Frozen Large lAnguage Models Enable data-efficient language-image pre-training) that leverages frozen large language models as text encoders, naturally processing long text inputs and demonstrating impressive multilingual generalization. FLAME comprises two key components: 1) a multifaceted prompt distillation technique for extracting diverse semantic representations from long captions, which better aligns with the multifaceted nature of images, and 2) a facet-decoupled attention mechanism, complemented by an offline embedding strategy, to ensure efficient computation. Extensive empirical evaluations demonstrate FLAME’s superior performance. When trained on CC3M, FLAME surpasses the previous state-of-the-art by 4.9% in ImageNet top-1 accuracy. On YFCC15M, FLAME surpasses the WIT-400M-trained CLIP by 44.4% in average image-to-text recall@1 across 36 languages, and by 34.6% in text-to-image recall@1 for long-context retrieval on Urban-1k. Code is available at https://github.com/MIV-XJTU/FLAME.
1 Introduction
Figure 1: Conceptual comparison of text streams. FLAME leverages frozen large language models (LLMs) to directly process long captions. With multifaceted prompts, this framework extracts diverse semantic embeddings, achieving data efficiency. Preserving LLMs’ inherent capabilities enables multilingual generalization.
Multimodal learning, particularly language-image pre-training, has made significant strides with models like CLIP (Contrastive Language-Image Pre-training) [35], which learn transferable visual representations through language supervision. CLIP-style models have demonstrated state-of-the-art performance across a wide range of downstream tasks [35, 51, 29, 39, 28, 45, 36]. However, the deployment of CLIP models in real-world scenarios presents several challenges, especially when data is scarce. Specifically, there are two major limitations that hinder their broader applicability: 1) Limited Training Pairs—high-quality image-text pairs, particularly those involving long-form descriptions and non-English languages, are difficult to come by; and 2) Constrained Model Capacity—the standard CLIP text encoder has a max sequence length limitation of 77 tokens, which restricts its ability to leverage the rich, long-context data that could be highly beneficial.
To address these challenges, recent works have primarily focused on data-centric solutions. Approaches such as synthesizing high-quality short captions using multimodal large language models [14, 24, 32] or translating short captions into multiple languages [43, 46, 8] attempt to augment training data. Some works also attempt to break down long captions into shorter segments [54] to fit the constraints of the CLIP text encoder. While these solutions are promising, they suffer from inherent limitations: The capacity bottleneck of the text encoder, particularly strict input restrictions, constrains the efficient utilization of data, and the naive decomposition strategy may lead to a loss of contextual meaning, resulting in suboptimal representations.
On the other hand, model-centric approaches—such as interpolating positional encodings to handle longer sequences [53]—attempt to extend the capabilities of CLIP’s text encoder. However, these solutions are fundamentally constrained by the design of the CLIP model, which allows a maximum of 77 tokens, with an effective token length of fewer than 20. These limitations often translate into suboptimal performance on downstream tasks [23]. Consequently, there remains a need for a more effective text encoder capable of processing long-form and multilingual content without relying on extensive data augmentation or architectural modifications.
In this paper, we propose a fundamentally new approach by utilizing frozen large language models (LLMs) as text encoders. Our key insight is that despite the conventional wisdom suggesting frozen text encoders lead to suboptimal performance [50, 14], powerful LLMs can provide sufficiently rich semantic representations for effective language-image pre-training. This strategy holds significant promise due to the inherent capabilities of LLMs: they are pre-trained on diverse multilingual corpora, excel at processing long-form text, and effectively extract semantic information at various granularities. While traditional decoder-only LLMs use causal attention mechanisms that limit their ability to generate rich contextual representations, recent advancements have shown that carefully designed prompting and attention mechanisms can enable effective text embedding for various tasks [21, 55, 52, 38, 6]. However, applying these techniques to language-image pre-training introduces unique challenges. Unlike general text embedding tasks where a single representation may suffice, visual content is inherently more complex, containing multifaceted semantic information that requires comprehensive understanding, especially when training data is limited.
To address these challenges, we introduce FLAME (Frozen Large lAnguage Models Enable data-efficient language-image pre-training), an innovative framework that leverages frozen LLMs to enhance language-image pre-training efficiency. FLAME employs a novel multifaceted prompt distillation technique to guide frozen LLMs in extracting various semantic facets from each image-text pair, maximizing the utility of limited training data. For example, when processing a long caption, FLAME simultaneously extracts semantic features ranging from fine-grained object attributes to abstract scene-level concepts. This approach facilitates a comprehensive visual-semantic alignment for each training sample, effectively enabling long-context understanding.
To ensure computational efficiency, FLAME incorporates a facet-decoupled attention mechanism that enables single-pass inference, preserving the independence of different semantic facets. This attention mechanism, combined with an offline embedding strategy, minimizes training overhead and improves the practical viability of our framework. By keeping the LLMs frozen, FLAME maintains the long-context and multilingual processing capabilities inherent in LLMs while reducing the need for expensive fine-tuning.
We evaluate FLAME across various data-scarce scenarios, demonstrating its superiority over previous methods. When trained on the CC3M dataset, FLAME outperforms the state-of-the-art model [54] by achieving a 4.9% improvement in ImageNet top-1 accuracy. In multilingual evaluation, FLAME surpasses WIT-400M-trained CLIP [35] by 44.4% in average image-to-text recall@1 across 36 languages on the Crossmodal-3600 dataset [40]. Moreover, FLAME excels in long-context image-text retrieval tasks, achieving a text-to-image recall@1 of 87.9% on the Urban-1k dataset [53], outperforming WIT-400M-trained CLIP by 34.6%. These results are particularly impressive considering that FLAME achieves these outcomes with a fraction of the training data used by previous methods.
Our key contributions are as follows:
  • We challenge the conventional wisdom about frozen text encoders and demonstrate that frozen LLMs can effectively enhance language-image pre-training through rich semantic representations.
  • FLAME introduces a novel framework that leverages frozen LLMs for data-efficient language-image pre-training via multifaceted prompt distillation.
  • We propose a facet-decoupled attention mechanism, complemented by an offline embedding strategy, to enhance computational efficiency while capturing comprehensive semantic representations, making the framework practical for real-world applications.
  • We demonstrate through extensive experimental validation that FLAME significantly outperforms existing methods in data-scarce scenarios, excelling in long-context understanding and multilingual tasks.
2 Related Works
Contrastive Language-Image Pre-training.
CLIP [35] pioneered transferable visual representation learning through language supervision, with subsequent works like ALIGN [19] scaling up through larger datasets. Recent efforts have focused on improving CLIP’s efficacy and data efficiency through various techniques: self-supervision [33, 30], modified losses [48, 15, 49, 25, 22], selective encoder tuning [50], and masked training [31, 12]. To enhance representation quality, researchers have explored hierarchical semantics [17], relaxed matching constraints [16], and caption synthesis using multimodal models [47, 14, 41, 24, 32, 54]. Efforts to handle long-form text [53] and multilingual content [46, 43, 8, 11] have relied on position encoding modifications and extensive data synthesis. In contrast to these approaches that work within CLIP’s inherent constraints, our framework fundamentally resolves both the length limitation and multilingual challenges by leveraging the rich capabilities of frozen LLMs.
Large Language Models as Text Encoders.
While large language models (LLMs) [7, 2, 42, 3, 20, 1, 4] excel at text understanding, their decoder-only architecture with causal attention poses challenges for generating rich text representations. Recent works have explored various ways to extract embeddings from LLMs, including contrastive training on paired data [44, 26, 34], single-word distillation [21, 55, 27, 52], sentence repetition [38], and bidirectional attention with modified training objectives [6]. Our work advances this direction by introducing multifaceted prompt distillation and facet-decoupled attention, specifically designed for language-image pre-training. Unlike previous methods that focus on general text embedding, our approach extracts comprehensive visual-semantic representations spanning multiple granularities while maintaining computational efficiency through innovative attention mechanisms.
3 Method
Language-image pre-training faces two fundamental challenges: limited availability of high-quality training pairs (especially for long-context and multilingual scenarios) and restricted token length of text encoders. While previous works address these challenges through data synthesis [14, 32, 24, 54] or architecture modifications [53], these solutions heavily remain constrained by CLIP’s fundamental design limitations.
We propose FLAME (Frozen Large lAnguage Models Enable data-efficient language-image pre-training), introducing a paradigm shift by leveraging frozen large language models (LLMs) as text encoders. This approach challenges the conventional wisdom that frozen text encoders lead to suboptimal performance [50, 14]. We demonstrate that powerful frozen LLMs, when properly utilized, can enable effective language-image pre-training through their rich semantic representations. FLAME inherits LLMs’ native capabilities in processing long-form text and understanding multiple languages, while enabling efficient training through optimized text embedding computation.
3.1 Framework Overview
Figure 2: FLAME overview. This framework harnesses the sophisticated long-text comprehension capabilities of large language models to conduct language-image pre-training directly on long captions. Based on multifaceted prompts, it extracts a diverse array of representations embedded within the long caption, thereby enhancing semantic alignment.
FLAME leverages frozen LLMs for visual-semantic alignment by exploiting their sophisticated text comprehension capabilities without any fine-tuning. As illustrated in Figure 2, our framework consists of three key components: 1) a frozen LLM that transforms each long caption into multiple complementary semantic representations through our proposed multifaceted prompt distillation technique, 2) a trainable visual encoder that distills the image content into a unified global representation, and 3) a learnable projection $\phi$ implemented as a two-layer multilayer perceptron (MLP) that aligns these heterogeneous embeddings in a shared semantic space for contrastive learning.
Given a batch of $N$ image-text pairs $\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents an image and $y_{i}$ is its corresponding long caption, we extract multiple semantic perspectives using a set of prompts $\{P_{k}\}_{k=1}^{K}$. The training objective is a symmetric contrastive loss: $\mathcal{L}=\frac{1}{2}(\mathcal{L}_{I}+\mathcal{L}_{T})$, where the text-to-image component $\mathcal{L}_{I}$ is formulated as:
$\small\mathcal{L}_{I}=-\sum_{i=1}^{N}\sum_{k=1}^{K}\log\frac{\exp(\cos\langle f_{t}(y_{i},P_{k}),\phi(f_{v}(x_{i}))\rangle/\tau)}{\sum_{j=1}^{N}\exp(\cos\langle f_{t}(y_{j},P_{k}),\phi(f_{v}(x_{i}))\rangle/\tau)},$ (1)
where $f_{t}(y_{i},P_{k})\in\mathbb{R}^{d_{t}}$ generates text embeddings through the LLM with prompt $P_{k}$, $f_{v}(x_{i})\in\mathbb{R}^{d_{v}}$ extracts visual features through the encoder, $\phi(\cdot)$ is the two-layer MLP that projects visual features into the text embedding space, and $\tau$ is a learnable temperature parameter. The image-to-text loss $\mathcal{L}_{T}$ follows the same formulation with image and text roles reversed.
The following sections detail our two key technical contributions: multifaceted prompt distillation for extracting diverse semantic representations from long captions (Section 3.2), and facet-decoupled attention for efficient computation (Section 3.3).
3.2 Multifaceted Prompt Distillation
While LLMs have shown promising capabilities in natural language understanding, their decoder-only architecture with causal attention inherently limits their ability to generate rich contextual representations. Previous works attempt to address this limitation either through bidirectional attention retraining [6] or by extracting the final token’s features [21, 55, 52]. However, these approaches only produce a single universal representation, which is insufficient for language-image pre-training where images contain multiple semantic facets from objects to scenes. To address this challenge, we introduce multifaceted prompt distillation, a systematic approach that decomposes visual semantics into a hierarchical structure mirroring human visual perception, from local object properties to global scene understanding.
Hierarchical Semantic Decomposition.
We identify three fundamental levels of visual semantics that are crucial for comprehensive image understanding:
  • Entity Level: Captures both primary and secondary objects in the scene, essential for object-centric understanding. This level is further divided into object categories and their distinctive attributes, enabling fine-grained visual discrimination.
  • Interaction Level: Focuses on dynamic elements such as actions and events, crucial for understanding the temporal and relational aspects of the scene. This bridges the gap between static object recognition and dynamic scene interpretation.
  • Scene Level: Abstracts high-level concepts such as atmosphere and emotion, capturing the holistic context that humans naturally perceive beyond individual objects and actions.
Figure 3: Facet-decoupled attention. By streamlining all prompts with a shared prefix and applying this facet-decoupled attention mask, the overhead of feature extraction is greatly reduced.
Prompt Engineering Principles.
Based on this hierarchical decomposition, we design our prompts following three key principles:
  • Semantic distinctiveness: Each prompt targets a distinct aspect of visual semantics to minimize redundancy in the extracted representations.
  • Constrained Output Space: All prompts are designed to generate single-word responses, forcing the LLM to distill complex visual concepts into concise, discriminative features.
  • Cognitive Alignment: Prompt structure mirrors human visual processing by explicitly separating different levels of semantic abstraction.
Our prompt set $\{P_{k}\}_{k=1}^{K}$ systematically covers the identified semantic hierarchy. Each prompt follows a template structure as shown below, where text in blue indicates the shared prefix containing the image caption $y_{i}$, and red marks the position where we extract the final hidden state feature from LLM:
Detailed image description: "$y_{i}$". After thinking step by step, the prominent characteristic or pattern of the main object in this image means in just one word:"
This hierarchical decomposition enables more robust visual-semantic alignment by matching images and text at multiple semantic levels simultaneously, while providing interpretable insights into the model’s understanding of visual content through its structured representation of semantic concepts. The complete list of our prompts is provided in the supplementary material.
3.3 Facet-Decoupled Attention
While our multifaceted approach provides rich semantic representations, a naive implementation would require $K$ separate forward passes through the LLM, resulting in significant computational overhead. To achieve efficient multi-perspective feature extraction, we propose an innovative attention mechanism design that enables single-pass inference while preserving semantic independence.
Our key insight is that all prompts share an identical prefix structure which dominates the input sequence length. Specifically, each prompt contains the core context: “Detailed image description: $y_{i}$. After thinking step by step,”. This observation motivates us to transform $K$ independent sequences into a single structured input:
$\{\langle y_{i},\mathrm{prefix},P_{k}\rangle\}_{k=1}^{K}\rightarrow\langle y_{i},\mathrm{prefix},\{P_{k}\}_{k=1}^{K}\rangle$
However, with this concatenated structure, standard causal attention would allow information leakage between different semantic facets, compromising the independence of extracted features. We address this challenge through facet-decoupled attention, which enforces strict boundaries between semantic perspectives using a specially designed attention mask (Figure 3). The mask is implemented through an efficient prefix KV Cache distribution strategy: by computing the prefix’s KV Cache once and distributing it for parallel prompt computation, our mechanism enables efficient extraction of multifaceted features while maintaining their semantic distinctiveness.
Notably, our use of a frozen LLM enables a significant practical advantage: we can pre-compute and cache all prompt-specific text embeddings offline. This design choice eliminates the LLM forward passes during training, resulting in competitive memory consumption and training speed compared to the original CLIP, while maintaining richer semantic representations. The computational overhead is thus reduced to a one-time preprocessing step, enabling efficient training at scale.
(a) Text Retrieval
(b) Image Retrieval
Figure 4: Multilingual zero-shot retrieval recall@1 results on Crossmodal-3600. Left: text retrieval. Right: image retrieval. Despite being trained solely on English datasets, FLAME achieves outstanding average performance across all 36 languages, surpassing mSigLIP, which is trained on the multilingual WebLI dataset with 100 languages.
4 Experiments
4.1 Datasets and Implementation Details
Dataset Details.
We use two primary datasets for language-image pre-training: CC3M [37] and YFCC15M [30]. Our ablation studies are primarily conducted using the CC3M dataset. Given the differences in the availability of web-crawled data, we re-implement the data synthesis pipeline from DreamLIP [54] using MiniCPM-Llama3-V 2.5111https://huggingface.co/openbmb/MiniCPM-Llama3-V-2_5 to generate both long and short captions. The raw captions are also included during training. All these datasets will be released to the community to benefit other research.
Implementation Details.
Our method is implemented using OpenCLIP [^2] https://github.com/mlfoundations/open_clip [18] and trained on a single node equipped with 8 NVIDIA A800 GPUs. The evaluation follows the CLIP benchmark protocol [^3] https://github.com/LAION-AI/CLIP_benchmark or the respective official benchmark protocols. We use ViT-B/16 [13, 5] as the default visual encoder and Mistral-Nemo [^4] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 [20] as the default large language model (LLM). For tasks with short text input, we apply a single prompt for inference, which is the first prompt at the scene level provided in the supplementary material. For a fair comparison with previous methods, we train all models for 32 epochs. Input images are resized to 224 $\times$ 224. The learnable temperature parameter $\tau$ is initialized to 0.07. More detailed hyper-parameters are provided in the supplementary material.
4.2 Zero-shot Image-Text Retrieval
{NiceTabular}
l—l—cc—cc & S4V-val Urban-1k
Method Dataset I2T T2I I2T T2I
CLIP [35] CC3M 21.4 20.2 10.7 9.6
Long-CLIP [53] CC3M+S4V 51.3 46.1 15.5 18.5
FLAME CC3M 85.6 80.0 65.3 66.6
CLIP [35] YFCC15M 57.1 45.9 30.4 23.6
Long-CLIP [53] YFCC15M+S4V 77.2 77.6 40.5 46.1
FLAME YFCC15M
94.1 93.2 84.0 87.9
CLIP [35] WIT-400M 78.2 79.6 67.5 53.3
Table 1: Long-context retrieval recall@1 results. S4V denotes ShareGPT4V. FLAME trained on YFCC15M exhibits a remarkable 34.6% improvement over WIT-400M-trained CLIP in text-to-image retrieval performance on Urban-1k.
{NiceTabular}
l—l—cccccc—cccccc Text Retrieval Image Retrieval
Dataset Method MSCOCO Flickr30k MSCOCO Flickr30k
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CC3M CLIP [35] 8.7 23.9 33.7 7.1 19.7 28.6 17.4 37.9 50.1 13.9 30.8 40.5
MLLM-A [32] 35.9 62.4 73.9 63.5 86.6 91.7 26.5 51.1 62.7 49.3 74.8 83.1
DreamLIP [54] 39.9 67.2 78.1 66.8 89.6 94.4 29.8 55.2 66.3 50.7 76.7 83.6
FLAME 43.3 69.1 78.9 67.3 87.6 93.1 28.6 54.5 65.7 53.6 79.9 87.1
YFCC15M CLIP [35] 30.7 56.2 67.4 54.9 80.0 88.4 19.1 40.9 52.5 37.2 64.3 74.3
SoftCLIP [16] 30.9 56.2 68.3 56.2 82.1 88.6 19.2 41.2 52.6 37.2 64.3 74.5
DreamLIP [54] 55.8 80.7 88.7 84.9
97.3 99.1 42.3 68.9 78.0 65.3 86.7 91.8
FLAME
60.5 82.9 89.3 86.4 97.3 98.6 43.9 70.4 79.7 73.3 91.7 95.5
WIT-400M CLIP [35] 52.4 76.7 84.7 81.9 96.2 98.8 33.1 58.4 69.0 62.1 85.6 91.8
Table 2: Short-context retrieval. The best results are in bold and the second best are underlined. Our FLAME outperforms CLIP while using a significantly smaller dataset for training (YFCC15M v.s. WIT-400M).
Long-Context Retrieval.
FLAME excels in long-context image-text retrieval tasks, as demonstrated by evaluations on ShareGPT4V-val [9] and Urban-1k [53]. Despite being trained on YFCC15M, FLAME achieves an 87.9% recall@1 in Urban-1k text-to-image retrieval, representing a 34.6% improvement over CLIP trained on the much larger WIT-400M dataset. Additionally, FLAME demonstrates an average improvement of 14.8% on ShareGPT4V-val. We also re-implement the recent method based on positional encoding interpolation [53], i.e., pre-training a CLIP model on CC3M and YFCC15M, then using their official code [^5] https://github.com/beichenzbc/Long-CLIP for fine-tuning on ShareGPT4V. The gap in results further demonstrates the advantage of FLAME’s paradigm shift over the interpolation-based architectural change on the standard CLIP text encoder.
Multilingual Retrieval.
FLAME leverages the intrinsic capabilities of pre-trained LLMs, enabling effective multilingual generalization. Although the model is trained on English datasets, it can be directly applied to downstream tasks in multiple languages without additional training. We evaluate this multilingual retrieval capability by conducting zero-shot retrieval across 36 languages from the Crossmodal-3600 dataset [40]. As shown in Figure 4, FLAME outperforms WIT-400M-trained CLIP [35] across all 36 languages. Specifically, FLAME achieves an average image-to-text recall@1 of 51.7% and text-to-image recall@1 of 49.0%, surpassing CLIP by margins of 44.4% and 44.5%, respectively. Notably, FLAME outperforms mSigLIP [51], which is trained on the multilingual WebLI [10] dataset with 100 languages, by 4.3% in image-to-text recall@1 and 14.6% in text-to-image recall@1.
{NiceTabular}
l—l—cccccccccc—c—c Dataset Method Food-101 CIFAR-10 CIFAR-100 SUN397 Cars Aircraft DTD Pets Caltech-101 Flowers Average ImageNet
CC3M CLIP [35] 10.6 53.9 20.4 31.2 1.2 1.1 10.4 11.7 43.2 12.9 19.7 16.0
LaCLIP [14] 14.2 57.1 27.5 35.1 1.6 1.6 16.6 15.6 52.7 14.7 23.7 21.5
MLLM-A [32] 18.7 58.4 32.4 43.8 3.9 1.5 20.2 32.1 63.5 17.5 29.2 25.0
DreamLIP [54] 19.4 74.3 44.2 45.9 2.8 1.0 17.0 27.1 63.1 14.7 31.0 31.1
FLAME 32.1 73.6 42.0 56.6 6.7 6.9 43.8 41.2 74.1 26.3 40.3 36.0
YFCC15M CLIP [35] 35.0 67.1 34.8 42.0 5.1 6.3 13.9 20.4 54.5 44.3 32.3 34.1
DreamLIP [54] 44.2
89.0 62.0 57.1 9.2 6.4 30.5 32.6 79.8 40.2 45.1 48.2
FLAME
61.8 86.1 56.7 66.8 10.7 10.3 54.9 40.7 78.9 51.7 51.9 51.5
Table 3: Zero-shot classification. FLAME significantly improves top-1 accuracy on ImageNet and average top-1 accuracy on 10 downstream datasets over the previous methods.
{NiceTabular}
l—l—cccccc Evaluation
Method Training ar en it jp zh avg
CLIP [35] en 0.4 68.3 21.8 4.2 1.4 19.2
CN-CLIP [46] en+zh 0.1 27.7 7.5 14.6 47.0 19.4
NLLB-CLIP [43] en+multi 18.4 24.3 18.0 16.5 16.6 18.8
FLAME en 32.7 51.5 42.1 36.9 39.7 40.6
Table 4: Multilingual ImageNet1k classification. CLIP is trained on WIT-400M. CN-CLIP is pre-trained on WIT-400M and fine-tuned on 200M Chinese data. NLLB-CLIP is pre-trained on WIT-400M and fine-tuned on LAION-COCO-NLLB with 200 languages. FLAME is trained on YFCC15M. Unlike other methods, FLAME generalizes well from monolingual training to multilingual evaluation.
Short-Context Retrieval.
FLAME also significantly outperforms CLIP in short-context retrieval tasks. When trained on YFCC15M, it shows an 8.7%/4.4% average improvement on MSCOCO/Flickr30k compared to CLIP trained on WIT-400M, as summarized in Table 4.2. These results indicate that while FLAME is designed for fine-grained semantic learning of long texts, it also achieves competent representations for context-limited texts, highlighting its potential in more general downstream scenarios.
4.3 Image Classification
Zero-Shot Classification.
We evaluate FLAME on zero-shot image classification tasks across diverse datasets. As shown in Table 4.2, FLAME achieves a 4.9% and 3.3% improvement in ImageNet top-1 accuracy when trained on CC3M and YFCC15M, respectively. On 10 common downstream datasets, FLAME achieves an average accuracy of 40.3%/51.9%, improving by 9.3%/6.8% over the prior state-of-the-art [54]. These results highlight that FLAME obtains powerful classification capabilities by aligning fine-grained entity-level representations.
Multilingual Classification.
To evaluate multilingual capabilities in image classification of FLAME, we conduct experiments on the multilingual ImageNet1k benchmark, covering Arabic (ar), English (en), Italian (it), Japanese (ja), and Chinese (zh). Results in Table 4.2 show that while CLIP trained on WIT-400M performs well in English, it lacks multilingual performance. We find that models like CN-CLIP [46] trained on specific languages compromise their performance in others. In contrast, FLAME generalizes effectively across all five languages, maintaining strong performance even when trained only on an English dataset. Additionally, our results comprehensively outperform models trained on multilingual datasets [43], with an average accuracy improvement of 21.8%.
Linear-Probe Classification.
We further evaluate FLAME’s feature extraction capabilities through linear-probe experiments, as shown in Table 4.3. On downstream datasets, FLAME improves by 3.4% and 3.0% on average when trained on CC3M and YFCC15M, respectively. These results demonstrate that aligning visual encoders with LLMs enhances generalization across downstream tasks.
{NiceTabular}
l—l—ccccccc—c Dataset Method Food-101 CIFAR-10 CIFAR-100 Cars Aircraft DTD Caltech-101 Average
CC3M CLIP [35] 62.6 86.8 68.1 32.8 40.9 63.4 82.0 62.4
LaCLIP [14] 63.8 87.7 69.5 32.4 42.7 64.0 83.3 63.3
MLLM-A [32] 64.0 87.7 68.5 34.5 32.1 60.4 85.5 61.8
DreamLIP [54] 71.2 92.2 74.0 31.5 26.7 70.4 88.5 64.9
FLAME 72.0 90.3 72.9 45.2 38.6 69.7 89.5 68.3
YFCC15M CLIP [35] 77.2 88.5 66.4 29.0 25.5 65.2 82.4 62.0
HiCLIP [17] 81.0 89.1 70.4 36.4 32.3 68.7 86.4 66.3
DreamLIP [54] 83.6
96.5 82.3 41.8 34.6 74.3 91.2 72.0
FLAME
85.9 95.0 81.0 54.3 39.3 76.8 92.5 75.0
Table 5: Linear-probe classification. FLAME achieves competent enhancements of average accuracy across downstream datasets.
4.4 Ablation Studies
{NiceTabular}
c—cc—cc—cc—cc—cc ShareGPT4V-val Urban-1k MSCOCO Flickr30k Classfication
# Prompts I2T@1 T2I@1 I2T@1 T2I@1 I2T@1 T2I@1 I2T@1 T2I@1 ImageNet DS Avg.
1 68.2 65.0 45.7 41.3 35.7 23.7 61.0 47.4 33.8 38.5
3 72.7 68.5 53.9 46.0 37.4 25.1 62.5 49.9 34.1 39.0
5 79.5 74.1 59.8 57.0 40.0 26.9 65.1 52.2 35.2 39.9
7 85.6 80.0 65.3 66.6 43.3 28.6 67.3 54.6 36.0 40.3
9 83.8
80.8 67.1 67.1 41.6 28.5 68.5 54.2 35.6 40.0
Table 6: Ablations on the number of prompts. Leveraging multifaceted prompts enhances performance, particularly in terms of long-context benchmarks. The right amount of prompts can achieve the best-balanced performance.
{NiceTabular}
cl—cc Backbone ImageNet DS Avg.
Text Mistral-7B 34.4 39.6
Mistral-Nemo 36.0 40.3
Llama-3.1-8B 34.1 39.8
Visual ViT-S/16 32.4 (+19.0) 37.1 (+20.7)
ViT-B/16 36.0 (+20.0) 40.3 (+20.6)
ViT-L/14 38.7 (+22.5)
42.9 (+22.2)
Table 7: Ablations on backbone architectures. The values in parentheses indicate improvements over the CLIP baseline using the same visual backbone. Whether employing larger or compact architectures, our framework demonstrates sufficient robustness.
Number of Prompts.
The quantity of prompts used plays a crucial role in the granularity of information extracted from long texts. To explore this relationship, we conducted an extensive ablation study on the number of prompts. The evaluation covers long/short-context retrieval and image classification tasks, with the results summarized in Table 4.4. DS refers to the downstream datasets from Table 4.2. We began with a diverse set of 9 prompts and performed random sampling to employ varying numbers of prompts for experimentation.
Our results show that using too few prompts leads to suboptimal performance, particularly on the long-context benchmarks. This phenomenon is due to the incomplete extraction and learning of features from long captions. Notably, the performance drop is less pronounced on classification tasks involving simpler short text inputs. On the other hand, when more than 7 prompts are used, the accuracy improvement plateaus. This suggests that beyond a certain point, the additional prompts provide diminishing returns, and the feature extraction from the long caption reaches a saturation point where further prompts may even hinder performance on short-context and classification tasks. Therefore, we use 7 as the default number of prompts to achieve a more balanced performance.
Different Backbone Architectures.
To assess the robustness and scalability of our framework, we conducted ablation studies using different backbone architectures. These included various text encoders (Mistral-7B [^6] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Mistral-Nemo, and Llama-3.1-8B [^7] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct) and visual encoders (ViT-S/16, ViT-B/16, and ViT-L/14). Whether reducing or increasing the architecture size, our framework continues to perform robustly. For instance, when using Mistral-7B, our model achieves an ImageNet top-1 accuracy of 34.4% and an average accuracy of 39.6% on downstream datasets, both of which significantly surpass the prior state-of-the-art results (31.1% and 31.0%, respectively) [54]. Additionally, we observe that performance improvements scale with the size of the visual encoder. For instance, FLAME with the ViT-L/14 encoder increases ImageNet top-1 accuracy by 22.5% over the CLIP baseline, while using ViT-S/16 yields a 19.0% improvement, underscoring the scalability of our approach.
Figure 5: Embedding and training overhead. FDA denotes facet-decoupled attention. While online training is hindered by inefficiencies, our approach of freezing LLMs allows for offline embedding to lower the overhead. Our facet-decoupled attention mechanism further boosts the efficiency of offline embedding, resulting in a 3.6x speedup.
Embedding and Training Overhead.
We evaluate the embedding and training overhead on CC3M, with results illustrated in Figure 5. For the recent method [54], we utilize their official code [^8] https://github.com/zyf0619sjtu/DreamLIP to measure training time. As shown, online training suffers from inefficiencies due to the frequent forward passes of LLMs. Our option of freezing LLMs, however, enables offline embedding preprocessing, significantly reducing unnecessary overhead. Furthermore, our proposed facet-decoupled attention greatly enhances the efficiency of offline embedding, achieving a 3.6x speedup. Our training time remains competitive with CLIP and outperforms that of the recent method [54].
5 Conclusion and Limitation
In this paper, we introduced FLAME, a novel language-image pre-training framework that harnesses frozen large language models as text encoders. Through multifaceted prompt distillation and facet-decoupled attention mechanisms, FLAME demonstrates superior performance across long-context understanding, multilingual generalization, and data-efficient visual-semantic alignment. While the framework introduces additional computational overhead during the initial text embedding extraction phase, our attention optimization achieves more than 3.6x speedup compared to naive implementation, making the preprocessing cost manageable and worthwhile given the significant performance gains. With the rapid advancement of more efficient LLMs and the framework’s demonstrated scalability across model sizes, FLAME represents a promising direction for future large-scale language-image pre-training.